# Enhancing Score-Based Sampling Methods with Ensembles

### v1

@@ Abstract
We introduce ensembles within score-based sampling methods to develop gradient-free approximate sampling techniques that leverage the collective dynamics of particle ensembles to compute approximate reverse diffusion drifts. We introduce the underlying methodology, emphasizing its relationship with generative diffusion models and the previously introduced Föllmer sampler. We demonstrate the efficacy of ensemble strategies through various examples, ranging from low- to medium-dimensionality sampling problems, including multi-modal and highly non-Gaussian probability distributions, and provide comparisons to traditional methods like NUTS. Our findings highlight the potential of ensemble strategies for modeling complex probability distributions in situations where gradients are unavailable. Finally, we showcase its application in the context of Bayesian inversion problems within the geophysical sciences.
Machine Learning, ICML
@@ 1 Introduction
Gradient-free methods for sampling from probability distributions are crucial in scenarios where gradient information is unavailable or computationally prohibitive. Unfortunately, many performant samplers such as the No U-Turn Sampler (NUTS) or the Metropolis Adjusted Langevin Algorithm (MALA) require gradient information to fully realize their advantages (Roberts & Tweedie, 1996; Hoffman et al., 2014). Others have investigated the use of transport based sampling methods to generate samples from probability distributions through stochastic differential equations, e.g., Huang et al. (2021); Vargas et al. (2023). In this context, Huang et al. (2021) have suggested a Monte Carlo approximation of the Föllmer drift to sample from probability distributions (Föllmer, 1985, 1986). They show that this approach results in an approximation to the solution of the Schrödinger bridge problem (Schrödinger, 1932). We build on these ideas and introduce ensemble strategies that leverage the collective dynamics of a particle ensemble to approximately compute the score function within a reverse diffusion process. Compared to the aforementioned Föllmer sampler, a goal is to reduce evaluations of the probability distribution that one wants to sample from. In the setting of Bayesian inverse problems, this helps to reduce forward model evaluations, thereby offering an efficient sampling technique. More concretely, we introduce an importance sampling Monte Carlo estimator for the score function of a forward diffusion process in order to sample from a probability distribution using the associated reverse diffusion process. Because there is some flexibility in the choice of forward diffusion process, we can design samplers that are still useful at medium- to high-dimensionality.
This work details the foundational methodology of our ensemble strategies, emphasizing its connection with generative diffusion models and the pre-existing Föllmer sampler. We assess the performance of our method through a range of diverse examples, encompassing low- to medium-dimensional Bayesian inference problems, including multi-modal and non-elliptical distributions. Our comparative analysis with Markov Chain Monte Carlo (MCMC) methods that utilize gradient information (e.g., Roberts & Tweedie, 1996; Hoffman et al., 2014) highlights the effectiveness of ensemble strategies in modeling complex probability distributions when gradients are not available. Furthermore, we demonstrate the practical application of our method in Bayesian inversion problems, particularly in the geophysical sciences where the dimensionality of the problem at hand can be of medium to large size, showcasing its potential for sophisticated probabilistic modeling.
@@ 1.1 Related Work
This paper builds upon existing research in stochastic processes, particularly focusing on the integration of Schrödinger-Föllmer approaches and particle and ensemble methods in computational physics (see, for example, Reich (2019) or Calvello et al. (2022) for recent overviews). Recent contributions adjacent to this area also include the work by Boffi & Vanden-Eijnden (2023) on the probability flow solution of the Fokker-Planck equation and the study by Bunne et al. (2022) on recovering stochastic dynamics through Gaussian Schrödinger bridges. Further, Dai et al. (2021) explored aspects of global optimization using Schrödinger-Föllmer diffusion. Huang et al. (2021) and Jiao et al. (2021) provided an initial analysis of the error and convergence properties of Schrödinger-Föllmer samplers in non-convex settings and introduced ideas this work was inspired by. Building on this, Zhang & Chen (2022) and Vargas et al. (2023) demonstrated the use of neural networks to improve the efficiency of Föllmer drift estimates in high-dimensional settings, and Bortoli et al. (2023) related the Schrödinger-Föllmer perspective to score-based generative models. In the area of ensemble-based gradient-free sampling techniques, Ensemble Kalman Inversion (EKI; e.g., Schillings & Stuart (2017), Ensemble Kalman Sampling (EKS; Garbuno-Inigo et al. (2020)), and related variants have been developed to address situation where gradients are unavailable. See Chada (2022) or the aforementioned Calvello et al. (2022) for overviews.
@@ 1.2 Our Contributions
This work introduces ensemble strategies to calculate estimates of score functions. Our main contribution is an importance sampling Monte Carlo estimator that reuses the samples being generated in a reverse diffusion process to define an importance sampling distribution. This approach has the advantage of keeping the Monte Carlo samples necessary for estimating the score function of the reverse diffusion process low. Furthermore, we conduct extensive experiments with a hierarchy of problems, comparing our method against MCMC methods that utilize gradient information to improve sample efficiency (e.g., NUTS by Hoffman et al. (2014)). Lastly, we explore various parameters influencing performance, such as noise schedules, importance sampling strategies, and ensemble sizes, providing insights into its application in complex systems modeling. We also show how flexibility resulting from the choice of forward diffusion process can lead to performant algorithms that efficiently leverage prior information.
2 Theory
This section introduces the theoretical foundations for what follows, establishing the mathematical framework and principles that guide our methodology. We first give a quick overview of score-based sampling with diffusion models and then address some of the strategies that make our methods work for the demonstrated experiments.
@@ 2.1 Forward & Reverse Diffusion Processes
where $\boldsymbol{b}_{t}$ is a time-dependent linear operator, $\boldsymbol{\mu}$ is the mean vector the process reverts to, and $\boldsymbol{g}_{t}$ is a time-dependent scaling matrix (often referred to as the drift and diffusion operators, respectively). The associated reverse diffusion process (Anderson, 1982) is given by
@@
where $\boldsymbol{s}_{t}$ is the score of the solution of the forward diffusion equation and where Equation (2) is integrated backwards in time. This forward-reverse duality has been leveraged in numerous studies within the areas of generative machine learning to generate samples from high dimensional data distributions, e.g., Song & Ermon (2019); Song et al. (2021) and follow-on studies. For a given initial probability distribution $p_{0}$ that we want to sample from, we can write down an exact expression for the score related to the process in Equation (1):
$\boldsymbol{s}_{t}(\boldsymbol{x})$ $=\nabla\log p_{t}(\boldsymbol{x})$ (3a)
$p_{t}(\boldsymbol{x})$ $=\int\kappa_{t}(\boldsymbol{x}|\boldsymbol{x}^{\prime})p_{0}(\boldsymbol{x}^{\prime})\,\mathrm{d}\boldsymbol{x}^{\prime}$ (3b)
@@
where $p_{t}$ is the solution to the Fokker-Planck equation associated with Equation (1) and the kernel function $\kappa_{t}$ is a Gaussian corresponding to the associated Green’s function. The kernel mean and covariance are given (Gardiner et al., 1985) as
$\boldsymbol{\mu}_{t}(\boldsymbol{x}^{\prime})=\left(1-e^{-\int_{0}^{t^{\prime}}\boldsymbol{b}_{t^{\prime}}\,\mathrm{d}t^{\prime}}\right)\boldsymbol{\mu}+e^{-\int_{0}^{t}\boldsymbol{b}_{t^{\prime}}\,\mathrm{d}t^{\prime}}\boldsymbol{x}^{\prime}$ (4a)
$\boldsymbol{\Sigma}_{t}=\int_{0}^{t}e^{-\int_{t^{\prime}}^{t}\boldsymbol{b}_{t^{\prime\prime}}\,\mathrm{d}t^{\prime\prime}}\boldsymbol{g}_{t^{\prime}}\boldsymbol{g}_{t^{\prime}}^{T}e^{-\int_{t^{\prime}}^{t}\boldsymbol{b}_{t^{\prime\prime}}^{T}\,\mathrm{d}t^{\prime\prime}}\,\mathrm{d}t^{\prime}.$ (4b)
Having an analytic kernel is possible because the forward diffusion process is of Ornstein-Uhlenbeck form. We can see that the solution to the forward diffusion equation interpolates between the initial distribution $p_{0}$ and a normal distribution as $t\rightarrow\infty$. In general, we choose $\boldsymbol{b}_{t}$ in such a manner that at $t=1$ the forward diffusion has approximately converged to this normal distribution.
2.2 Bayesian Context
In a Bayesian context, when $p_{0}$ is a posterior distribution, and if the prior distribution is a normal distribution with a given mean and covariance, then this approach can be used to approximately transport samples from the prior to the posterior over a time span $[0,1]$ by choosing $\boldsymbol{b}_{t}$ and $\boldsymbol{g}_{t}$ appropriately. Given the kernel of the forward diffusion process, one can use the reverse diffusion equation to draw samples from $p_{0}$. However, the main challenge lies in the fact that the score $\boldsymbol{s}_{t}$ is only available as a potentially high-dimensional integral. In the next sections, we introduce an approach that uses an ensemble of samples to define an importance sampling estimator for the score, thereby alleviating this problem for low- to medium-dimensional problems.
@@ 2.3 Estimating the Score Function
A key aspect of our ensemble strategies is the estimation of the score within the reverse diffusion process, which is essential for accurately simulating particle trajectories in Equation (2). This section outlines our approach to estimating the score function. To that end, we introduce an importance sampling distribution $p_{\rm{is}}$ (e.g., Elvira & Martino, 2022) such that we can rewrite the solution to the forward diffusion equation as
@@
where $N$ is the number of Monte Carlo samples drawn. Here, our goal is to define $p_{\rm{is}}$ in such a way that the reverse diffusion process generates good samples from the posterior distribution $p_{0}$ even though we use an approximation for the score. When actually computing the estimate $\hat{\boldsymbol{s}}_{t}(\boldsymbol{x})=\nabla\log\hat{p}_{t}(\boldsymbol{x})$ for the score function, we follow the approach of Vargas et al. (2023) and use the logsumexp trick to dramatically improve the numerical stability of the score estimation. We further discuss several variance reduction strategies for the Monte Carlo estimator $\hat{p}_{t}$ in the next section.
@@ 2.4 Variance Reduction Strategies
This section focuses on the implementation of various variance reduction strategies that are essential for enhancing the accuracy and efficiency of the sample generation process.
@@ 2.4.1 Importance Sampling via Sample Ensemble
$\boldsymbol{\mu}_{\rm{ens}}$ $=\frac{\sum_{i=1}^{N_{\rm{ens}}}\boldsymbol{x}_{i}}{N_{\rm{ens}}}$ (7a)
$\boldsymbol{\Sigma}_{\rm{ens}}$ $=\frac{\sum_{i=1}^{N_{\rm{ens}}}\left(\boldsymbol{x}_{i}-\boldsymbol{\mu}_{\rm{ens}}\right)\left(\boldsymbol{x}_{i}-\boldsymbol{\mu}_{\rm{ens}}\right)^{T}}{N_{\rm{ens}}}$ (7b)
@@
Utilizing this ensemble-based $p_{\rm{is}}$ therefore relies on the adequacy of the normal distribution. In practice, we also set the number of Monte Carlos samples $N$ equal to the ensemble size $N_{\rm{ens}}$. We outline an alternative importance sampling distribution using a multiple importance sampling strategy based on the ensemble of samples in Appendix E. We find that both distributions work similarly well for the examples presented in this paper. The actual variance reduction achieved is directly related to the ensemble size; that is, as the ensemble size tends to infinity, the score estimator becomes exact. In practice, we find that we do not need to actually draw samples from $p_{\rm{is}}$ when calculating estimators for $\hat{p}_{t}$ or $\hat{\boldsymbol{s}}_{t}$. Instead, we find that evaluating the Monte Carlo approximation at the ensemble members gives equally reasonable results.

### v2

# Gradient-Free Score-Based Sampling Methods with Ensembles

@@ Abstract
Recent developments in generative modeling have utilized score-based methods coupled with stochastic differential equations to sample from complex probability distributions. However, these and other performant sampling methods generally require gradients of the target probability distribution, which can be unavailable or computationally prohibitive in many scientific and engineering applications. Here, we introduce ensembles within score-based sampling methods to develop gradient-free approximate sampling techniques that leverage the collective dynamics of particle ensembles to compute approximate reverse diffusion drifts. We introduce the underlying methodology, emphasizing its relationship with generative diffusion models and the previously introduced Föllmer sampler. We demonstrate the efficacy of the ensemble strategies through various examples, ranging from low- to medium-dimensionality sampling problems, including multi-modal and highly non-Gaussian probability distributions, and provide comparisons to traditional methods like the No-U-Turn Sampler. Additionally, we showcase these strategies in the context of a high-dimensional Bayesian inversion problem within the geophysical sciences. Our findings highlight the potential of ensemble strategies for modeling complex probability distributions in situations where gradients are unavailable.
keywords:
Uncertainty quantification , Brownian motion , Bayesian inference , Sampling theory , Gradient-free sampling , Geophysical parameter estimation
journal: Applied Mathematical Modelling
[…]
@@ 1 Introduction
Gradient-free methods for sampling from probability distributions are necessary in scenarios where gradient information is unavailable or computationally prohibitive. Such scenarios are common in Bayesian inverse problems involving complex, black-box simulators, legacy scientific code, or models with inherently non-differentiable components. Unfortunately, many performant samplers such as the No U-Turn Sampler (NUTS) or the Metropolis Adjusted Langevin Algorithm (MALA) require gradient information to fully realize their advantages [1, 2]. Recent investigations have utilized transport-based sampling methods to generate samples from probability distributions through stochastic differential equations (SDEs), e.g., [3, 4]. Diffusion models, developed within the field of generative modeling, can also be viewed as solving SDEs linking unknown data distributions to known prior distributions [5]. In this context, Huang et al. [3] suggested a Monte Carlo approximation of the Föllmer drift to sample from probability distributions [6, 7], showing that this approach results in an approximation to the solution of the Schrödinger bridge problem.
We build on these ideas and introduce ensemble strategies that leverage the collective dynamics of a particle ensemble to approximate the score function (gradient of the log probability density function with respect to data) within a reverse diffusion process without actually needing to compute gradients. We develop these strategies with the aim to sample from arbitrary probability distributions known only up to a normalization constant, which is common in Bayesian inference applications. Compared to the aforementioned Föllmer sampler, our goal is to reduce the number of evaluations of the target probability distribution. In the setting of Bayesian inverse problems, this goal helps to reduce forward model evaluations, thereby offering an efficient sampling technique. More concretely, we introduce an importance sampling Monte Carlo estimator for the score function of a forward diffusion process in order to sample from a probability distribution using the associated reverse diffusion process. Moreover, we exploit the flexibility in the choice of the forward diffusion process to design samplers that are useful at medium- to high-dimensionality.
This work details the foundational methodology of our ensemble strategies, emphasizing its connection with generative diffusion models and the pre-existing Föllmer sampler. We assess the performance of our method through a range of diverse examples, encompassing low- to medium-dimensional Bayesian inference problems, including multi-modal and non-elliptical distributions. Our comparative analysis with Markov Chain Monte Carlo (MCMC) methods that utilize gradient information (e.g., [1, 2]) highlights the effectiveness of ensemble strategies in modeling complex probability distributions when gradients are not available. Furthermore, we demonstrate the practical application of our method in Bayesian inversion problems, particularly in the geophysical sciences where the dimensionality of the problem at hand can be of medium to large size, showcasing its potential for sophisticated probabilistic modeling.
@@ 1.1 Related Work
This paper builds upon existing research in stochastic processes, particularly focusing on the integration of Schrödinger-Föllmer approaches and particle and ensemble methods in computational physics (see, for example, [8] or [9] for recent overviews). Recent contributions adjacent to this area also include the work by Boffi and Vanden-Eijnden [10] on the probability flow solution of the Fokker-Planck equation and the study by Bunne et al. [11] on recovering stochastic dynamics through Gaussian Schrödinger bridges. Further, Dai et al. [12] explored aspects of global optimization using Schrödinger-Föllmer diffusion. Huang et al. [3] and Jiao et al. [13] provided an initial analysis of the error and convergence properties of Schrödinger-Föllmer samplers in non-convex settings and introduced ideas this work was inspired by. Building on this, Zhang and Chen [14] and Vargas et al. [4] demonstrated the use of neural networks to improve the efficiency of Föllmer drift estimates in high-dimensional settings, and Bortoli et al. [15] related the Schrödinger-Föllmer perspective to score-based generative models. In the area of ensemble-based gradient-free sampling techniques, Ensemble Kalman Inversion (EKI; [16]), Ensemble Kalman Sampling (EKS; [17]), and related variants have been developed to address situation where gradients are unavailable. See Chada [18] or the aforementioned Calvello et al. [9] for overviews. Furthermore, ensemble based adaptive importance sampling techniques, such as the ensemble transport methods developed by Cotter et al. [19], share conceptual similarities with our approach by utilizing the statistics of the current particle ensemble to dynamically construct more effective importance sampling distributions or transport steps.
@@ 1.2 Our Contributions
This work introduces ensemble strategies to calculate estimates of score functions. Our main contribution is an importance sampling Monte Carlo estimator that reuses the samples being generated in a reverse diffusion process to define an importance sampling distribution. This approach has the advantage of keeping the number of Monte Carlo samples necessary for estimating the score function of the reverse diffusion process low. Furthermore, we conduct extensive experiments with a hierarchy of problems, comparing our method against MCMC methods that utilize gradient information to improve sample efficiency (e.g., NUTS by Hoffman and Gelman [2]). Lastly, we explore various parameters influencing performance, such as noise schedules, importance sampling strategies, and ensemble sizes, providing insights into its application in complex systems modeling. We also show how flexibility resulting from the choice of forward diffusion process can lead to performant algorithms that efficiently leverage prior information.
2 Foundations of Score-based Sampling with Diffusion Models
This section introduces the theoretical foundations for what follows, establishing the mathematical framework and principles that guide our methodology. We first provide a brief overview of score-based sampling with diffusion models and then address some of the strategies that make our methods work for the demonstrated experiments.
@@ 2.1 Forward & Reverse Diffusion Processes
where $\boldsymbol{b}_{t}$ is a time-dependent linear operator, $\boldsymbol{\mu}$ is the mean vector the process reverts to, and $\boldsymbol{g}_{t}$ is a time-dependent scaling matrix (often referred to as the drift and diffusion operators, respectively). The forward diffusion can be viewed as the incremental addition of noise to a data distribution. The associated reverse diffusion process [20] is given by
@@
where $\boldsymbol{s}_{t}$ is the time-dependent score of the solution of the forward diffusion equation and where Equation (2) is integrated backwards in time. Thus, once the score function is known for all $t$, samples can be generated from the original data distribution using the reverse diffusion SDE. Importantly, the score function allows for sampling from un-normalized probability distributions, which is imperative for applications like Bayesian inference [21]. This forward-reverse duality has been leveraged in numerous studies within the areas of generative machine learning to generate samples from high dimensional data distributions, e.g., [5, 22] and follow-on studies. As we demonstrate later with numerical examples, the diffusion framework provides flexibility in the construction of the forward and reverse processes (via selection of $\boldsymbol{b}_{t}$, $\boldsymbol{\mu}$, and $\boldsymbol{g}_{t}$), allowing for tuning of the sampling process for a given probability distribution.
For a given initial probability distribution $p_{0}$ that we want to sample from, we can write down an exact expression for the score related to the process in Equation (1):
$\boldsymbol{s}_{t}(\boldsymbol{x})$ $=\nabla_{\boldsymbol{x}}\log p_{t}(\boldsymbol{x}),$ (3a)
$p_{t}(\boldsymbol{x})$ $=\int\kappa_{t}(\boldsymbol{x}|\boldsymbol{x}^{\prime})p_{0}(\boldsymbol{x}^{\prime})\,\mathrm{d}\boldsymbol{x}^{\prime},$ (3b)
@@
where $p_{t}$ is the solution to the Fokker-Planck equation associated with Equation (1) and the kernel function $\kappa_{t}$ is a Gaussian corresponding to the associated transition probability (i.e., Green’s function) of the Ornstein-Uhlenbeck process. The kernel mean and covariance can be derived analytically as [23]
$\boldsymbol{\mu}_{t}(\boldsymbol{x}^{\prime})=\left(1-\exp\left(-\int_{0}^{t^{\prime}}\boldsymbol{b}_{t^{\prime}}\,\mathrm{d}t^{\prime}\right)\right)\boldsymbol{\mu}+\exp\left(-\int_{0}^{t}\boldsymbol{b}_{t^{\prime}}\,\mathrm{d}t^{\prime}\right)\boldsymbol{x}^{\prime},$ (4a)
$\boldsymbol{\Sigma}_{t}=\int_{0}^{t}\exp\left(-\int_{t^{\prime}}^{t}\boldsymbol{b}_{t^{\prime\prime}}\,\mathrm{d}t^{\prime\prime}\right)\boldsymbol{g}_{t^{\prime}}\boldsymbol{g}_{t^{\prime}}^{T}\exp\left(-\int_{t^{\prime}}^{t}\boldsymbol{b}_{t^{\prime\prime}}^{T}\,\mathrm{d}t^{\prime\prime}\right)\,\mathrm{d}t^{\prime}.$ (4b)
Now consider a 1D Ornstein-Uhlenbeck process defined by $\mu=0$, $g_{t}=\sigma$, and $b_{t}=b$ is constant in time. In this case, $\mu_{t}(x^{\prime})=x^{\prime}\exp(-t)$ and $\Sigma_{t}=(\sigma^{2}/(2b))(1-\exp(-2bt))$ such that as $t\rightarrow\infty$, $\mu_{t}\rightarrow 0$ and $\Sigma_{t}\rightarrow\sigma^{2}/(2b)$. In other words, the solution to the forward diffusion equation interpolates between the initial distribution $p_{0}$ and a normal distribution with zero mean and a fixed variance. For standard Brownian motion, $\mu_{t}(x^{\prime})=x^{\prime}$ and $\Sigma_{t}=\sigma^{2}t$, i.e. the forward diffusion approaches a normal distribution with infinite variance as $t\rightarrow\infty$.
2.2 Score-based Sampling in Bayesian Inference
In a Bayesian context, when $p_{0}$ is a posterior distribution, and if the prior distribution is a normal distribution with a given mean and covariance, then this approach can be used to approximately transport samples from the prior to the posterior over a finite time span $0\leq t\leq 1$ by choosing $\boldsymbol{b}_{t}$, $\boldsymbol{\mu}$, and $\boldsymbol{g}_{t}$ appropriately. Given the kernel of the forward diffusion process, one can then use the reverse diffusion equation to draw samples from $p_{0}$. This behavior can be used to tune the selection of the forward process parameters for a given sampling problem. However, the main challenge lies in the fact that the score $\boldsymbol{s}_{t}$ is only available as a potentially high-dimensional integral. In the next sections, we introduce an approach that uses an ensemble of samples to define an importance sampling estimator for the score, thereby alleviating this problem for low- to medium-dimensional problems.
3 Ensemble Strategies for Gradient-Free Score Estimation and Sampling
Building upon the diffusion framework, this section introduces ensemble-based strategies designed to estimate the score function and enable efficient sampling from target distributions without requiring gradient information. We detail the core Monte Carlo score estimation and various techniques to enhance its accuracy and applicability, particularly for complex and high-dimensional problems.
@@ 3.1 Estimating the Score Function
A key aspect of our ensemble strategies is the estimation of the score within the reverse diffusion process, which is essential for accurately simulating particle trajectories in Equation (2). This section outlines our approach to estimating the score function. To that end, we introduce an importance sampling distribution $p_{\rm{is}}$ (e.g., [24]) such that we can rewrite the solution to the forward diffusion equation as
@@
where $N$ is the number of Monte Carlo samples drawn. Here, our goal is to define $p_{\rm{is}}$ in such a way that the reverse diffusion process generates good samples from the posterior distribution $p_{0}$ even though we use an approximation for the score. When actually computing the estimate $\hat{\boldsymbol{s}}_{t}(\boldsymbol{x})=\nabla_{\boldsymbol{x}}\log\hat{p}_{t}(\boldsymbol{x})$ for the score function, we follow the approach of Vargas et al. [4] and use the logsumexp trick to dramatically improve the numerical stability of the score estimation. A flowchart summarizing the ensemble score-based sampling procedure (and variance reduction strategies for the Monte Carlo estimator $\hat{p}_{t}$ discussed in the next section) is shown in Figure 1.
Figure 1: Flowchart illustrating the core reverse diffusion sampling process and the ensemble-based estimation framework used to replace the score function when gradients cannot be computed, highlighting the roles of the different variance reduction and efficiency strategies.
@@ 3.2 Variance Reduction Strategies
This section focuses on the implementation of various variance reduction strategies that are essential for enhancing the accuracy and efficiency of the sample generation process. These strategies can be grouped into: i) general variance reduction strategies that we apply to all sampling problems (Sections 3.2.1-3.2.3); and ii) problem-specific strategies that utilize additional information to improve sample quality (Section 3.2.4).
@@ 3.2.1 Importance Sampling via Sample Ensemble
$\boldsymbol{\mu}_{\rm{ens}}$ $=\frac{\sum_{i=1}^{N_{\rm{ens}}}\boldsymbol{x}_{i}}{N_{\rm{ens}}},$ (7a)
$\boldsymbol{\Sigma}_{\rm{ens}}$ $=\frac{\sum_{i=1}^{N_{\rm{ens}}}\left(\boldsymbol{x}_{i}-\boldsymbol{\mu}_{\rm{ens}}\right)\left(\boldsymbol{x}_{i}-\boldsymbol{\mu}_{\rm{ens}}\right)^{T}}{N_{\rm{ens}}},$ (7b)
@@
Utilizing this ensemble-based $p_{\rm{is}}$ therefore relies on the adequacy of the normal distribution as an importance sampling distribution. In practice, we also set the number of Monte Carlos samples $N$ equal to the ensemble size $N_{\rm{ens}}$. We outline an alternative importance sampling distribution using a multiple importance sampling strategy based on the ensemble of samples in B. We find that both distributions work similarly well for the examples presented in this paper. The actual variance reduction achieved is directly related to the ensemble size; that is, as the ensemble size tends to infinity, the score estimator becomes exact (see C for a convergence analysis). In practice, we find that we do not need to actually draw samples from $p_{\rm{is}}$ when calculating estimators for $\hat{p}_{t}$ or $\hat{\boldsymbol{s}}_{t}$. Instead, we find that evaluating the Monte Carlo approximation at the ensemble members gives equally reasonable results.